• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Feuerbach

    Log entry time 16:10:33 on June 15, 2005

    Entry number 146181

    keyword=Plan for this evening

    Repairs on the L-HRS Q2 and Q3 are on going, but a set of broken components have been found. However, they do not expect to be finished until during tomorrow day shift.

    So tonight, the Hall will be swept and locked up at 2100. At this point, Eugene Chudakov will use the Moller polarimeter to make a measurement of the beam polarization, partially in service to Accelerator and partially to commission the device.

    During owl shift, Nilanga Liyanage will be given control to perform measure the effects of shielding on detector count rates.

    Once the tests are complete, inform MCC that we are finished, in case there are plans/tests that they wish to perform.

    Regardless of the status of these test, beam delivery will end in the morning in time for the hall to be surveyed at 6am and brought to Restricted Access, such that the repairs on Q3 can continue.

    If there are any questions of what to do, please contact the Run Coordinator at 867-1787.